OpenStudy (anonymous):

pressure p and volume v of a gas are related by the law p(v^u)=k where u and k are constants.Find the rate of pressure with time.

OpenStudy (anonymous):

dp/dt= -k/(v^2u)

OpenStudy (anonymous):

how?

OpenStudy (ranga):

Does your original equation look like this? \[pv ^{u} = k\]

OpenStudy (anonymous):

@ranga ,yes!

OpenStudy (ranga):

First multiply both sides by: \[v ^{-u}\]

OpenStudy (ranga):

Then find dp/dt

OpenStudy (anonymous):

i used the rules Btw i'm nt sure about it

OpenStudy (anonymous):

But if i multiply throughout by v^-u im left with; \[p=k v ^{-u}\] P is still not varying with 't' so how can i find dp/dt?

OpenStudy (anonymous):

@Ahmad1919 ,thanks for giving it a try,buddy!

OpenStudy (anonymous):

if f(x) = a/b^k f'(x) = -a/b^2k it's a rule and ur wlc

OpenStudy (anonymous):

oh! thanks!

OpenStudy (anonymous):

*yet again,haha

OpenStudy (ranga):

@kesh, \[p = kv ^{-u}\] \[\frac{ dp }{ dt } = k(-u)v ^{-u-1}\frac{ dv }{ dt}\] \[\frac{ dp }{ dt} = \frac{ -ku }{ v ^{u+1}}\frac{ dv }{ dt }\]

OpenStudy (anonymous):

Thank you so much @ranga !